### v1

@@ I Introduction
Recently, we introduced a promising method Kang, Trizio, and Parrinello (2024); Trizio, Kang, and Parrinello (2025) based on the calculation of the committor function combined with the On-the-fly Probability Enhanced Sampling (OPES) method. Invernizzi and Parrinello (2020); Trizio et al. (2024) This approach has demonstrated unique capabilities in uniformly sampling entire reaction pathways and has several merits: it is based on a physics-grounded variational principle, it enables exhaustive sampling of the transition state (TS) region, and it provides a powerful tool to reveal the reaction process in remarkable detail. In the method section, we summarize the key concepts of this approach, while a more extensive discussion can be found in the original papers. Kang, Trizio, and Parrinello (2024); Trizio, Kang, and Parrinello (2025)
The central quantity of this approach is the committor function $q(\mathbf{x})$, which, for the transition between two states $A$ and $B$, gives the probability that a configuration $\mathbf{x}$ evolves to state $B$ before having first passed by AAWeinan and Vanden-Eijnden (2010) Owing to this probabilistic description, the committor is widely considered to provide the ideal description of a reactive process, and due to its many attractive features, much work has been done to determine it. Bolhuis, Dellago, and Chandler (2000); Ma and Dinner (2005); Berezhkovskii and Szabo (2005); Peters and Trout (2006); E. and Vanden-Eijnden (2006); Li, Lin, and Ren (2019); Khoo, Lu, and Ying (2019); Rotskoff, Mitchell, and Vanden-Eijnden (2022); He, Chipot, and Roux (2022); Chen et al. (2023); Chen, Roux, and Chipot (2023); Jung et al. (2023); Mitchell and Rotskoff (2024) However, computing $q(\mathbf{x})$ is challenging, as it requires extensive sampling of the TS region. Traditionally, this was obtained either from long simulations Krivov (2021) or via transition path sampling. Bolhuis et al. (2002); Bolhuis and Csányi (2018); Dellago, Bolhuis, and Geissler (2006); Jung et al. (2023) Recently, we have shown that once the $A$ and $B$ states are known, the committor can be computed via a self-consistent iterative procedure. Kang, Trizio, and Parrinello (2024); Trizio, Kang, and Parrinello (2025) Our approach is based on a variational principle first pioneered by OnsagerOnsager (1938) and later framed into the formalism of Kolmogorov’s backward equation. Kolmogoroff (1931) This principle states that, under the assumption of overdamped Langevin dynamics, the committor can be determined by imposing the boundary conditions $q(\mathbf{x}_{i\in A})=0$ and $q(\mathbf{x}_{i\in B})=1$ and minimizing the functional
@@
In Ref. 5, the committor was expressed as a feed-forward neural network $q_{\boldsymbol{\theta}}(\mathbf{d}(\mathbf{x}))$, where $\boldsymbol{\theta}$ are learnable parameters. The input of the neural network is a set of $N_{d}$ physical descriptors $\mathbf{d}(\mathbf{x})$ which provide a convenient way to enforce the system’s physical symmetries and favor interpretability. The model parameters are optimized using as a loss function the $K[q(\mathbf{x})]$ functional supplemented by the above-mentioned boundary conditions.
Due to its ability to extensively and uniformly sample the reactive process, our approach has proven robust and successful in applications that range from biophysics Das et al. ; Berselli et al. (2026); Trizio, Kang, and Parrinello (2025) to chemistry and materials science. Trizio, Kang, and Parrinello (2025); Kang, Trizio, and Parrinello (2024); Deng et al. (2026) However, in some cases, training the committor can be rather expensive. The main numerical source of inefficiency is often the calculation of the gradients relative to the atomic coordinates in the variational functional. The cost of computing these derivatives can be rather high when using complex descriptors that depend on a large number of atoms. Deng et al. (2026); Dietrich et al. (2024)
@@ II Methods
In the following, we start by giving an overview of the whole iterative procedure in Sec. II.1, highlighting the key differences with Ref. 5. Then, we discuss in detail all the components of such a procedure. Specifically, in Sec. II.2, we recall the general features of the committor function and the main concepts of the original approach of Ref. 5. Then, in Sec. II.3, we present in detail the modifications to the variational loss function used for the optimization of the approximated committor models.
@@
Here, we outline the steps of the self-consistent iterative procedure proposed here for the learning of approximated committor functions. Such a procedure closely resembles the original recipe of Ref.5, except for the use of a different loss function that leads to an approximated committor model, as detailed in the following sections.
@@ II.2 Background
Kolmogorov bias. To overcome the sampling limitations associated with the TS in rare events, in Ref. 4, we introduced the Kolmogorov bias potential $V_{K}(\mathbf{x})$ to stabilize the TS region and favor its sampling. This TS-oriented bias is defined as a functional of the committor, exploiting the strong localization of its gradients on the TS
@@
In practice, when applying $V_{K}(\mathbf{x})$ in enhanced sampling calculations, one can equivalently compute it considering the gradients with respect to the atomic coordinates or to the input descriptors, as discussed in Ref. 4. As the second possibility is simpler to implement and computationally cheaper, it is to be preferred.
@@
which offers a practical way to identify the transition state ensemble (TSE). Kang, Trizio, and Parrinello (2024); Trizio, Kang, and Parrinello (2025); Das et al. ; Berselli et al. (2026)
Machine learning the committor. In Ref. 5, we leveraged the variational principle of Eq. 1 to optimize a machine-learning-based parametrization of the committor through a self-consistent iterative procedure. During such a procedure, cycles of (enhanced) sampling and training are alternated to progressively improve both the committor estimate and the sampling until convergence is reached.
@@ II.3 Reformulation of variational loss
While the procedure summarized above provides a reliable framework that has already been successfully applied to a range of systems, Kang, Trizio, and Parrinello (2024); Trizio, Kang, and Parrinello (2025); Das et al. ; Deng et al. (2026); Berselli et al. (2026); Kang et al. (2026) it still presents some practical drawbacks in certain cases. Indeed, during training, the main computational cost arises from the minimization of the variational term, which requires computing the gradients $\nabla_{\mathbf{u}}q_{\boldsymbol{\theta}}(\mathbf{x})$ with respect to the mass-scaled atomic positions $\mathbf{u}$, which in many applications may become particularly demanding. Specifically, using automatic differentiation packages, Paszke et al. (2019) the computational load of the training is proportional to i) the number of descriptors used, ii) their complexity, and iii) the number of atoms involved in the calculation of such descriptors. Even if this is not a conceptual limitation and strategies to reduce the overload can be devised, in practice, this can introduce an extra layer of complexity. This motivated us to seek a reformulation of our learning approach that would reduce the computational burden and simplify the set up protocol.
@@
which, after simple manipulations, fully detailed in the SI, can be written as
@@
As a final practical note, we recall that, for improved numerical stability, it is also possible to minimize $\log(L)$, as proposed in Ref. 30. In addition, we found that the new learning objective is more prone to overfitting in some cases. It is thus advisable to monitor this behavior and, if necessary, adopt strategies to avoid it, e.g., via early stopping or shorter training.
@@ III.1 Alanine Dipeptide conformational equilibrium
The first test system for our methodology was the conformational equilibrium of alanine dipeptide, often used as a benchmark for new enhanced sampling methods (see Fig. 1A). As often done in the past for this molecule, we use the interatomic distances between the heavy atoms of the molecule as input descriptors. Bonati, Rizzi, and Parrinello (2020); Trizio and Parrinello (2021); Trizio, Kang, and Parrinello (2025) Based on this description of the system, we set up our simplified approach, obtaining an approximated committor function, which was used to drive enhanced sampling calculations with the OPES and Kolmogorov biases. This provided effective sampling of the entire relevant phase space by promoting frequent transitions between the metastable states (see Fig. 1B). From the data thus collected, we recovered accurate free energy estimates that were almost indistinguishable from those obtained using as CVs the conventional $\varphi$-$\psi$ torsional angles or a committor model trained on the original variational approach (see Fig. 1C).
@@ III.3 OAMe-G2 binding
As a prototypical example of this class of processes, we studied the binding of the G2 molecule to the OAMe calixarene guest from the SAMPL5 challenge, Yin et al. (2017) which has been extensively characterized in several papers Bhakat and Söderhjelm (2017); Rizzi et al. (2021); Ray, Trizio, and Parrinello (2023); Trizio, Kang, and Parrinello (2025) (see Fig. 3A). As done in Refs. 38 and 5, we chose to treat the solvation degrees of freedom separately from the trivial positional ones. To this end, we trained our model solely on water coordination numbers, computed for a few ligand atoms and a set of virtual atoms above the binding pocket, and biased the resulting committor-like CV using both OPES and $V_{K}(\mathbf{x})$. Specifically, the OPES component of the bias was applied to both the committor-based CV $z(\mathbf{x})$ and the ligand-pocket distance. Following this strategy, we were able to promote efficient sampling and obtain accurate free energy estimates in good agreement with previous studies with a similar computational setup (see Fig. 3B and C).
@@
Figure 3: G2-OAMe binding. A. Representative snapshots of the binding stages, with the host molecule colored in pink and the guest in orange. Water molecules are represented as blue spheres in solid color when inside the pocket, and transparent otherwise. B. Free energy surface (FES) plotted along the learned $z$ CV. C. Convergence with simulation time of the estimated binding energy. The dashed line shows the reference value obtained with the same setup as Ref. 38, with the $\pm 0.5\,$kBT interval marked by the thin dotted lines. In panels B and C, average values obtained from 3 independent simulations are depicted as solid lines, whereas the corresponding standard deviations are shown as shaded areas.
@@ III.4 Crystallization of Si
Indeed, following Ref. 41, we employed as input descriptors the peaks of the anisotropic structure factor of silicon. In this case, considering the different relative magnitudes of these inputs, we also introduced a preprocessing normalization layer to the NN architecture. In addition, to make up for the expected larger free energy barrier and the considerably high temperature, we used a larger $\lambda$ value, i.e., $\lambda=1000$.
Even starting from data limited to the pure phases, our iterative procedure allowed us to progressively learn an effective reaction coordinate, which correlates well with the commonly used, but rather expensive, CVs based on environment similarity Piaggi and Parrinello (2019) and Steinhardt parameters (see SI, Fig. S7).
@@ IV Discussion
This paper fits into a broader effort to extend enhanced sampling methods to the detailed study of transition states and to make them more automated and accessible. In Refs. 4 and 5, we leveraged machine learning together with the longstanding formal theory of the committor function to achieve first extensive sampling of the TS region, and later uniform and ergodic sampling of the relevant phase space. With Ref. 32, we moved a step further toward automation by exploiting the expressive power of graph neural networks to operate directly on atomic coordinates, thereby bypassing the need for manually designed input descriptors, although at a higher computational cost.
@@ Acknowledgements.
The authors are grateful to Luigi Bonati, Florian Dietrich, Ioannis Galdadas, and Alice Triveri for useful feedback and discussions on the manuscript.
@@
Simulations details The alanine dipeptide (Ace-Ala-Nme) simulations in vacuum were carried out targeting the NVT ensemble using the GROMACS-2024.5 Abraham et al. (2015) MD engine patched with PLUMED-2.9.2 Tribello et al. (2014); consortium (2019) and the Amber99-SB Salomon-Ferrer, Case, and Walker (2013) force field with a 2 fs timestep. The Langevin dynamics is sampled with damping coefficient $\gamma_{i}=\frac{m_{i}}{\tau-t}$ with $\tau-t=0.05$ ps and target temperature of 300 K.
@@
Simulations details The intramolecular proton transfer simulations were carried out targeting the NVT ensemble using the CP2K-8.1 Kühne et al. (2020) software package patched with PLUMED-2.9.1 consortium (2019) at PM6 semi-empirical level. The integration step was 0.5 fs and we used the velocity rescaling thermostat Bussi, Donadio, and Parrinello (2007) set at 300 K with a time constant of 100 fs.
@@
Simulations details We performed the simulations with GROMACS-2024.5 Abraham et al. (2015) in combination with the PLUMED-2.9.4 consortium (2019) plugin. We use the GAFF Wang et al. (2004) force field with RESP Bayly et al. (1993) charges and the TIP3P Jorgensen et al. (1983) water model. We set the timestep to 2 fs, targeting the NVT ensemble at 300 K via a velocity rescale thermostat Bussi, Donadio, and Parrinello (2007) with a time constant of 0.1 ps. The simulation box is cubic with a side of 40.27 $\mathrm{\SIUnitSymbolAngstrom}$ and it contains 2100 water molecules in solution together with the host OAMe and the G2 guest molecule. Sodium ions are included to counterbalance excess charges. At every simulation step, the coordinates are aligned so that the vertical axis of the box coincides with the binding axis $h$, and the simulation box is centered on the virtual atom V1.
@@
Approximated committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the same water coordination numbers used in Ref. 38, 8 with respect to 2.5 Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)) as inputs of a neural network (NN) with architecture [12, 24, 8, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$. The training was performed for 5000 epochs. The $\alpha$ hyperparameter in the loss function was progressively diminished through the iterations, i.e., 103, 101, 10-3, and we optimized the $\log$ of the total loss for numerical stability. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations with the corresponding model are summarized in Table S3 alongside the simulation time $t_{s}$ and the output sampling time $t_{o}$.
@@
Simulations details Silicon crystallization simulations were carried out targeting the NPT ensemble using the same setup of Ref.Bonati, Piccini, and Parrinello (2021), employing LAMMPS Thompson et al. (2022) (22Jul2025 update1) patched with PLUMED-2.9.0. consortium (2019) As an interatomic potential we used the Stillinger–Weber interatomic potential. Stillinger and Weber (1985). A 3 × 3 × 3 supercell (216 atoms) was simulated in the isothermal–isobaric (NPT) ensemble with a time step of 2 fs. The velocity-rescaling thermostat Bussi, Donadio, and Parrinello (2007) with a target temperature of 1700 K and a relaxation time of 100 fs was used, while the barostat Martyna, Tobias, and Klein (1994) target pressure was set to 1 atm and the relaxation time to 1 ps. First, two 2 ns-long simulations of standard MD in the solid and liquid states were performed. During the simulation, we monitored the fraction of diamond-like atoms computed in PLUMED consortium (2019) with the Environment Similarity CV, Piaggi and Parrinello (2019) with parameters SIGMA = 0.4, LATTICE_CONSTANTS = 5.43, MORE_THAN = $\{$ R_0 = 0.5 NN = 12 MM = 24}\}.

### v2

@@ I Introduction
Recently, we introduced a promising method Kang et al. (2024); Trizio et al. (2025) based on the calculation of the committor function combined with the On-the-fly Probability Enhanced Sampling (OPES) method. Invernizzi and Parrinello (2020); Trizio et al. (2024) This approach has demonstrated unique capabilities in uniformly sampling entire reaction pathways and has several merits: it is based on a physics-grounded variational principle, it enables exhaustive sampling of the transition state (TS) region, and it provides a powerful tool to reveal the reaction process in remarkable detail. In the method section, we summarize the key concepts of this approach, while a more extensive discussion can be found in the original papers. Kang et al. (2024); Trizio et al. (2025)
The central quantity of this approach is the committor function $q(\mathbf{x})$, which, for the transition between two states $A$ and $B$, gives the probability that a configuration $\mathbf{x}$ evolves to state $B$ before having first passed by AAWeinan and Vanden-Eijnden (2010) Owing to this probabilistic description, the committor is widely considered to provide the ideal description of a reactive process, and due to its many attractive features, much work has been done to determine it. Bolhuis et al. (2000); Ma and Dinner (2005); Berezhkovskii and Szabo (2005); Peters and Trout (2006); E. and Vanden-Eijnden (2006); Li et al. (2019); Khoo et al. (2019); Rotskoff et al. (2022); He et al. (2022); Chen et al. (2023b, a); Jung et al. (2023); Mitchell and Rotskoff (2024) However, computing $q(\mathbf{x})$ is challenging, as it requires extensive sampling of the TS region. Traditionally, this was obtained either from long simulations Krivov (2021) or via transition path sampling. Bolhuis et al. (2002); Bolhuis and Csányi (2018); Dellago et al. (2006); Jung et al. (2023) Recently, we have shown that once the $A$ and $B$ states are known, the committor can be computed via a self-consistent iterative procedure. Kang et al. (2024); Trizio et al. (2025) Our approach is based on a variational principle first pioneered by OnsagerOnsager (1938) and later framed into the formalism of Kolmogorov’s backward equation. Kolmogoroff (1931) This principle states that, under the assumption of overdamped Langevin dynamics, the committor can be determined by imposing the boundary conditions $q(\mathbf{x}_{i\in A})=0$ and $q(\mathbf{x}_{i\in B})=1$ and minimizing the functional
@@
In Ref. 39, the committor was expressed as a feed-forward neural network $q_{\boldsymbol{\theta}}(\mathbf{d}(\mathbf{x}))$, where $\boldsymbol{\theta}$ are learnable parameters. The input of the neural network is a set of $N_{d}$ physical descriptors $\mathbf{d}(\mathbf{x})$ which provide a convenient way to enforce the system’s physical symmetries and favor interpretability. The model parameters are optimized using as a loss function the $K[q(\mathbf{x})]$ functional supplemented by the above-mentioned boundary conditions.
Due to its ability to extensively and uniformly sample the reactive process, our approach has proven robust and successful in applications that range from biophysics Das et al. ; Berselli et al. (2026); Trizio et al. (2025) to chemistry and materials science. Trizio et al. (2025); Kang et al. (2024); Deng et al. (2026) However, in some cases, training the committor can be rather expensive. The main numerical source of inefficiency is often the calculation of the gradients relative to the atomic coordinates in the variational functional. The cost of computing these derivatives can be rather high when using complex descriptors that depend on a large number of atoms. Deng et al. (2026); Dietrich et al. (2024)
@@ II Methods
In the following, we start by giving an overview of the whole iterative procedure in Sec. II.1, highlighting the key differences with Ref. 39. Then, we discuss in detail all the components of such a procedure. Specifically, in Sec. II.2, we recall the general features of the committor function and the main concepts of the original approach of Ref. 39. Then, in Sec. II.3, we present in detail the modifications to the variational loss function used for the optimization of the approximated committor models.
@@
Here, we outline the steps of the self-consistent iterative procedure proposed here for the learning of approximated committor functions. Such a procedure closely resembles the original recipe of Ref.39, except for the use of a different loss function that leads to an approximated committor model, as detailed in the following sections.
@@ II.2 Background
Kolmogorov bias. To overcome the sampling limitations associated with the TS in rare events, in Ref. 23, we introduced the Kolmogorov bias potential $V_{K}(\mathbf{x})$ to stabilize the TS region and favor its sampling. This TS-oriented bias is defined as a functional of the committor, exploiting the strong localization of its gradients on the TS
@@
In practice, when applying $V_{K}(\mathbf{x})$ in enhanced sampling calculations, one can equivalently compute it considering the gradients with respect to the atomic coordinates or to the input descriptors, as discussed in Ref. 23. As the second possibility is simpler to implement and computationally cheaper, it is to be preferred.
@@
which offers a practical way to identify the transition state ensemble (TSE). Kang et al. (2024); Trizio et al. (2025); Das et al. ; Berselli et al. (2026)
Machine learning the committor. In Ref. 39, we leveraged the variational principle of Eq. 1 to optimize a machine-learning-based parametrization of the committor through a self-consistent iterative procedure. During such a procedure, cycles of (enhanced) sampling and training are alternated to progressively improve both the committor estimate and the sampling until convergence is reached.
@@ II.3 Reformulation of variational loss
While the procedure summarized above provides a reliable framework that has already been successfully applied to a range of systems, Kang et al. (2024); Trizio et al. (2025); Das et al. ; Deng et al. (2026); Berselli et al. (2026); Kang et al. (2026) it still presents some practical drawbacks in certain cases. Indeed, during training, the main computational cost arises from the minimization of the variational term, which requires computing the gradients $\nabla_{\mathbf{u}}q_{\boldsymbol{\theta}}(\mathbf{x})$ with respect to the mass-scaled atomic positions $\mathbf{u}$, which in many applications may become particularly demanding. Specifically, using automatic differentiation packages, Paszke et al. (2019) the computational load of the training is proportional to i) the number of descriptors used, ii) their complexity, and iii) the number of atoms involved in the calculation of such descriptors. Even if this is not a conceptual limitation and strategies to reduce the overload can be devised, in practice, this can introduce an extra layer of complexity. This motivated us to seek a reformulation of our learning approach that would reduce the computational burden and simplify the set up protocol.
@@
which, after simple manipulations, fully detailed in the supplementary material, can be written as
@@
As a final practical note, we recall that, for improved numerical stability, it is also possible to minimize $\log(L)$, as proposed in Ref. 15. In addition, we found that the new learning objective is more prone to overfitting in some cases. It is thus advisable to monitor this behavior and, if necessary, adopt strategies to avoid it, e.g., via early stopping or shorter training.
@@ III.1 Alanine Dipeptide conformational equilibrium
The first test system for our methodology was the conformational equilibrium of alanine dipeptide, often used as a benchmark for new enhanced sampling methods (see Fig. 1A). As often done in the past for this molecule, we use the interatomic distances between the heavy atoms of the molecule as input descriptors. Bonati et al. (2020, 2021); Trizio and Parrinello (2021); Trizio et al. (2025) Based on this description of the system, we set up our simplified approach, obtaining an approximated committor function, which was used to drive enhanced sampling calculations with the OPES and Kolmogorov biases. This provided effective sampling of the entire relevant phase space by promoting frequent transitions between the metastable states (see Fig. 1B). From the data thus collected, we recovered accurate free energy estimates that were almost indistinguishable from those obtained using as CVs the conventional $\varphi$-$\psi$ torsional angles or a committor model trained on the original variational approach (see Fig. 1C).
@@ III.3 OAMe-G2 binding
As a prototypical example of this class of processes, we studied the binding of the G2 molecule to the OAMe calixarene guest from the SAMPL5 challenge, Yin et al. (2017) which has been extensively characterized in several papers Bhakat and Söderhjelm (2017); Rizzi et al. (2021); Ray et al. (2023); Trizio et al. (2025) (see Fig. 3A). As done in Refs. 37 and 39, we chose to treat the solvation degrees of freedom separately from the trivial positional ones. To this end, we trained our model solely on water coordination numbers, computed for a few ligand atoms and a set of virtual atoms above the binding pocket, and biased the resulting committor-like CV using both OPES and $V_{K}(\mathbf{x})$. Specifically, the OPES component of the bias was applied to both the committor-based CV $z(\mathbf{x})$ and the ligand-pocket distance. Following this strategy, we were able to promote efficient sampling and obtain accurate free energy estimates in good agreement with previous studies with a similar computational setup (see Fig. 3B and C).
@@
Figure 3: G2-OAMe binding. A. Representative snapshots of the binding stages, with the host molecule colored in pink and the guest in orange. Water molecules are represented as blue spheres in solid color when inside the pocket, and transparent otherwise. B. Free energy surface (FES) plotted along the learned $z$ CV. C. Convergence with simulation time of the estimated binding energy. The dashed line shows the reference value obtained with the same setup as Ref. 37, with the $\pm 0.5\,$kBT interval marked by the thin dotted lines. In panels B and C, average values obtained from 3 independent simulations are depicted as solid lines, whereas the corresponding standard deviations are shown as shaded areas.
@@ III.4 Crystallization of Si
Indeed, following Ref. 7, we employed as input descriptors the peaks of the anisotropic structure factor of silicon. In this case, considering the different relative magnitudes of these inputs, we also introduced a preprocessing normalization layer to the NN architecture. In addition, to make up for the expected larger free energy barrier and the considerably high temperature, we used a larger $\lambda$ value, i.e., $\lambda=1000$.
Even starting from data limited to the pure phases, our iterative procedure allowed us to progressively learn an effective reaction coordinate, which correlates well with the commonly used, but rather expensive, CVs based on environment similarity Piaggi and Parrinello (2019) and Steinhardt parameters (see supplementary material, Fig. S7).
@@ IV Discussion
This paper fits into a broader effort to extend enhanced sampling methods to the detailed study of transition states and to make them more automated and accessible. In Refs. 23 and 39, we leveraged machine learning together with the longstanding formal theory of the committor function to achieve first extensive sampling of the TS region, and later uniform and ergodic sampling of the relevant phase space. With Ref. 24, we moved a step further toward automation by exploiting the expressive power of graph neural networks to operate directly on atomic coordinates, thereby bypassing the need for manually designed input descriptors, although at a higher computational cost.
@@
Supplementary Material
The supplementary material includes the extended derivation of the simplified variational functional proposed here, additional figures, and the computational details of the simulations performed and training of the approximated committor models.
@@ Acknowledgements.
The authors are grateful to Luigi Bonati, Florian Dietrich, Ioannis Galdadas, Peilin Kang and Alice Triveri for useful feedback and discussions on the manuscript.
@@
Simulations details The alanine dipeptide (Ace-Ala-Nme) simulations in vacuum were carried out targeting the NVT ensemble using the GROMACS-2024.5 Abraham et al. (2015) MD engine patched with PLUMED-2.9.2 Tribello et al. (2014); consortium (2019) and the Amber99-SB Salomon-Ferrer et al. (2013) force field with a 2 fs timestep. The Langevin dynamics is sampled with damping coefficient $\gamma_{i}=\frac{m_{i}}{\tau-t}$ with $\tau-t=0.05$ ps and target temperature of 300 K.
@@
Simulations details The intramolecular proton transfer simulations were carried out targeting the NVT ensemble using the CP2K-8.1 Kühne et al. (2020) software package patched with PLUMED-2.9.1 consortium (2019) at PM6 semi-empirical level. The integration step was 0.5 fs and we used the velocity rescaling thermostat Bussi et al. (2007) set at 300 K with a time constant of 100 fs.
@@
Simulations details We performed the simulations with GROMACS-2024.5 Abraham et al. (2015) in combination with the PLUMED-2.9.4 consortium (2019) plugin. We use the GAFF Wang et al. (2004) force field with RESP Bayly et al. (1993) charges and the TIP3P Jorgensen et al. (1983) water model. We set the timestep to 2 fs, targeting the NVT ensemble at 300 K via a velocity rescale thermostat Bussi et al. (2007) with a time constant of 0.1 ps. The simulation box is cubic with a side of 40.27 $\mathrm{\SIUnitSymbolAngstrom}$ and it contains 2100 water molecules in solution together with the host OAMe and the G2 guest molecule. Sodium ions are included to counterbalance excess charges. At every simulation step, the coordinates are aligned so that the vertical axis of the box coincides with the binding axis $h$, and the simulation box is centered on the virtual atom V1.
@@
Approximated committor model training details To model the committor function $q_{\theta}(\mathbf{x})$ at each iteration, we used the same water coordination numbers used in Ref. 13, 8 with respect to 2.5 Å-spaced virtual atoms (V) along the binding axis and 4 with respect to 4 atoms on the ligand molecule (M) (2 on the ring (1 and 2), 2 on the terminal atoms (3 and 4)) as inputs of a neural network (NN) with architecture [12, 24, 8, 1] nodes/layer. For the optimization, we used the ADAM optimizer with an initial learning rate of $10^{-3}$ modulated by an exponential decay with multiplicative factor $\gamma=0.99995$. The training was performed for 5000 epochs. The $\alpha$ hyperparameter in the loss function was progressively diminished through the iterations, i.e., 103, 101, 10-3, and we optimized the $\log$ of the total loss for numerical stability. The number of iterations, the corresponding dataset size, and the $\lambda$ and the OPES BARRIER used in the biased simulations with the corresponding model are summarized in Table S3 alongside the simulation time $t_{s}$ and the output sampling time $t_{o}$.
@@
Simulations details Silicon crystallization simulations were carried out targeting the NPT ensemble using the same setup of Ref.Bonati et al. (2021), employing LAMMPS Thompson et al. (2022) (22Jul2025 update1) patched with PLUMED-2.9.0. consortium (2019) As an interatomic potential we used the Stillinger–Weber interatomic potential. Stillinger and Weber (1985). A 3 × 3 × 3 supercell (216 atoms) was simulated in the isothermal–isobaric (NPT) ensemble with a time step of 2 fs. The velocity-rescaling thermostat Bussi et al. (2007) with a target temperature of 1700 K and a relaxation time of 100 fs was used, while the barostat Martyna et al. (1994) target pressure was set to 1 atm and the relaxation time to 1 ps. First, two 2 ns-long simulations of standard MD in the solid and liquid states were performed. During the simulation, we monitored the fraction of diamond-like atoms computed in PLUMED consortium (2019) with the Environment Similarity CV, Piaggi and Parrinello (2019) with parameters SIGMA = 0.4, LATTICE_CONSTANTS = 5.43, MORE_THAN = $\{$ R_0 = 0.5 NN = 12 MM = 24}\}.